OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (10, -3) with a slope of -3/2. A. -13 B. 4 C. 13 D. -12

OpenStudy (anonymous):

\[slope =\frac{ -3 }{ 2 } =\frac{ 6-(-3) }{ x-(10) }\] now you need to solve for x \[x =2\frac{ 6+3 }{ -3 }+10=4\]